OpenStudy (hawaii_love):

solve for x 2x-20=70-x

OpenStudy (hawaii_love):

yes

alones (alones):

Add \(20\) to each side \(x=70−x+20 \)

alones (alones):

Than simplify \(70−x+20\) to \(90−x\) and add \(x\) : \(2x+x=90 \)

OpenStudy (steve816):

Than, as AloneS is saying, do some algebra\[x=30\]\[3x=90\]

OpenStudy (steve816):

Oops, I put that backwards

alones (alones):

Lol yess, but still so from there divide \(90\) by\(3\) \[\frac{ 90 }{ 3 }=?\]

alones (alones):

Yeah i got \(30\).

OpenStudy (steve816):